OpenStudy (anonymous):

HELP PLEASE MEDAL AWARDED TO ANYONE WHO DOES THIS CORRECTLY!!!! i need to make a system of linear equations with the y=mx + b equation and i don't know how to do it. I have the number but i just don't know where to put them or how to do it the numbers are 14.40 for m and 0 for b in the first one and the second is 11.99 for m and 0 b. First person to do correctly gets a medal and a big thanks!!

OpenStudy (anonymous):

y=mx +b (1) 14.4x + 0 = y (2) 11.99x + 0 = y Solve this and then u'll get the value of x

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

wait a minute the x is 3 though becuase there are three items being bought

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

3 items being bought......? You didn't post the original question so it's really hard to know what you're talking about.